Given r(x)=7x-22 and t(x)=x^2+3, express the sum (r+t)(x).
Answer by Edwin McCravy(20056)


Given r(x)=7x-22 and t(x)=x²+3, express the sum (r+t)(x).

Just add the right sides:

(r+t)(x) = r(x) + t(x) = (7x-22) + (x²+3) = 7x - 22 + x² + 3 = x² + 7x - 19

Edwin